# Distributed Multiple Fault Detection and Estimation in DC Microgrids with Unknown Power Loads

### v1

@@ Abstract.
This paper proposes a distributed diagnosis scheme to detect and estimate actuator and power line faults in DC microgrids subject to unknown power loads and stochastic noise. To address actuator faults, we design a fault estimation filter whose parameters are determined through a tractable optimization problem to achieve fault estimation, decoupling from power line faults, and robustness against noise. In contrast, the estimation of power line faults poses greater challenges due to the inherent coupling between fault currents and unknown power loads, which becomes ill-posed when the underlying system is insufficiently excited. To the best of our knowledge, this is the first study to address this critical yet underexplored issue. Our solution introduces a novel differentiate-before-estimate strategy. A set of diagnostic rules based on the temporal characteristics of a constructed residual is developed to distinguish load changes from line faults. Once a power line fault is detected, a regularized least-squares method is activated to estimate the fault currents, for which we further derive an upper bound on the estimation error. Finally, comprehensive simulation results validate the effectiveness of the proposed methods.
@@
Jingwei Dong, Per Mattsson, and André Teixeira are with the Division of Systems and Control, Uppsala University, Sweden ({jingwei.dong, per.mattson,andre.teixeira}@it.uu.se). Mahdieh S. Sadabadi is with the Department of Electrical and Electronic Engineering, The University of Manchester, Manchester, U.K. (e-mail: mahdieh.sadabadi@manchester.ac.uk).
@@ 1.1. Literature review
The existing results on fault diagnosis of DC microgrids typically focus on fault detection and isolation. The most straightforward way is to set detection thresholds based on characteristics of measured currents and voltages, as presented in [3]. While thresholding-based methods are easy to implement, their diagnosis performance is highly sensitive to the selection of thresholds, which necessitates a thorough understanding of signal properties in both normal and faulty conditions. By exploring additional time-domain information from measured signals, methods based on differential currents [4, 5] and traveling wave analysis [6] can achieve more accurate diagnosis results compared to simple thresholding-based approaches. For instance, the authors in [5] analyzed the first and second derivatives of fault currents to detect DC cable ground faults in the presence of load and operating mode changes. However, such a derivative-based approach is sensitive to noise and may lead to false alarms without a proper selection of thresholds.
In addition to time-domain information, frequency-domain methods, such as fast Fourier transform [7] and wavelet transform [8], are also effective for fault analysis in microgrid systems. Yet, their reliance on case-specific fault frequency characteristics limits their generalization capability. In recent years, the availability of large amounts of historical data has also facilitated the adoption of learning-based approaches in DC microgrid fault diagnosis. A key challenge, however, is the limited amount of data in fault scenarios. We refer interested readers to [9] and [10] for further insights.
@@
Let us first briefly review some observer-based fault diagnosis methods developed for DC microgrids. In [13], an unknown input observer was employed to detect arc faults on power lines. To enhance robustness against disturbances while maintaining fault sensitivity, the authors in [14] utilized the mixed $\mathcal{H}_{\_}/\mathcal{H}_{\infty}$ index in the design of an optimal fault detection observer for DC microgrids. Fault isolation can be achieved by building a bank of observers, each of which is designed to be sensitive to specific faults. For instance, the authors in [15] constructed fault matrices (or signatures) for different types of faults in DC microgrids, and then designed a bank of unknown input observers based on corresponding sensitivity matrices to distinguish fault types. In [16], a bank of unscented Kalman filters was developed to estimate system states under faults, whose outputs were then integrated to locate faulty sensors in DC microgrids. Regarding DAE-based approaches, they offer several advantages over conventional observer-based methods, including: (i) characterizing all possible residual generators for systems described by DAEs, and (ii) deriving residual generators of the lowest possible order [12, 17]. These benefits have also led to the development of anomaly detection methods for microgrid systems within the DAE framework, such as attack detection in [18] and ground fault detection in [19].
@@
Note that the aforementioned methods primarily focus on fault detection and isolation, while research on fault estimation for DC microgrids remains limited. Fault estimation involves determining the shape and size of faults, which has more stringent requirements than fault detection and isolation. The complexity is further exacerbated in DC microgrids with unknown power loads due to the introduction of nonlinearity [20]. To address this issue, the authors in [20] employed a sliding mode observer to estimate actuator and sensor faults in DC microgrids. More recently, an adaptive fault estimation method was developed in [21] to estimate the system states and faults simultaneously. However, both [20] and [21] rely on power line or load current measurements, which require additional sensors that increase system complexity and maintenance costs. The authors in [22] avoided this limitation by proposing a method based on an augmented state observer and an open-loop line current estimator, but the method is confined to addressing sensor attacks.
@@
Based on the above analysis, the problem of multiple fault estimation for DC microgrids remains insufficiently explored, particularly in scenarios where power line currents are unmeasurable and power loads are unknown. Our investigation finds that the coupled effects of fault-induced line currents and unknown power loads pose great challenges in distinguishing between them. This problem is even challenging when the system is not sufficiently excited (e.g., in the case of incipient faults), where the slow variation of signals leads to an ill-posed estimation problem. To the best of our knowledge, this is the first study to address this challenge in the context of DC microgrids.
@@ 1.2. Main contributions
  • Multiple fault estimation framework. We develop a multiple fault estimation framework for DC microgrids, which accounts for both actuator and power line faults in the presence of measurement noise, unknown power loads, and parametric uncertainties. This framework is built on fault estimation filters developed within the DAE framework, which enables suppression of noise and uncertainties, decoupling of disturbances, and estimation of multiple faults. The filter design with these requirements is then formulated as tractable optimization problems (Proposition 3.1).
  • Differentiation strategy for power line faults and step power load changes. Building on the practical assumption of piecewise-constant power loads, we propose a novel strategy to distinguish between power line faults and step changes in power loads by exploiting temporal characteristic differences of their effects on residuals (Proposition 4.1). This differentiation process is further formalized into a set of diagnosis rules (Equation (29)), facilitating the subsequent faulty line current estimation.
  • Faulty line current estimation with an explicit error bound. To address the ill-posed issue in faulty line current estimation, we formulate a regularized least-squares optimization method. An analytical solution is further derived to enable real-time estimation (Proposition 4.2). Furthermore, we establish a computable bound on the expectation of the estimation error, explicitly characterizing its dependence on load and fault variations (Theorem 4.3).
@@ 2.1. DG Dynamics
where $V_{i}(t)\in{\mathbb{R}}$ and $I_{t,i}(t)\in{\mathbb{R}}$ are the voltage at the Point of Common Coupling (PCC) $i$ and the filter current, respectively. The unknown power demand is denoted by $P_{i}(t)\in{\mathbb{R}}_{+}$. The control signal $u_{i}(t)=\varphi_{i}(t)V_{dc,i}$, where $\varphi_{i}(t)$ is the duty cycle of the buck converter. The current of power line $k$ is represented by $I_{k}(t)\in{\mathbb{R}}$ where $k\in\mathbb{N}_{L}$. The microgrid topology is encoded in the incidence matrix element $\mathbb{B}_{ik}$, which indicates the connection between power line $k$ and DG unit $i$. In particular, $\mathbb{B}_{ik}=1$ if DG $i$ is the positive end of line $k$, $\mathbb{B}_{ik}=-1$ if DG $i$ is the negative end of line $k$, and $\mathbb{B}_{ik}=0$ otherwise. Therefore, the term $\sum_{k=1}^{m}\mathbb{B}_{ik}I_{k}(t)$ is the total current injected into DG $i$ from its neighbors and represents the physical couplings of DG $i$ with neighboring DG units. The additive actuator fault in DG unit $i$ is denoted by $f_{a,i}(t)\in{\mathbb{R}}$.
@@ 2.2. Power Line Dynamics
where $R_{k}$ is the line resistance and $L_{k}$ is the line inductance. The fault on line $k$ is denoted by $f_{L,k}(t)\in{\mathbb{R}}$.
@@ 2.3. Stabilizing Voltage Controller
where $V_{i}^{*}$ is a reference voltage provided by a higher-level controller, $K_{i}=[k_{1,i}\ k_{2,i}\ k_{3,i}]$ is the voltage control gain vector. Design methods for $K_{i}$ can be found in [24] or [25].
@@ 2.4. Closed-loop Dynamics
where $f_{I,{i}}(t)=\sum_{k=1}^{m}\mathbb{B}_{ik}I_{k,f}(t)$ denote the aggregated faulty line current affecting DG unit $i$. The system matrices in (4) are as follows:
$\begin{split}A_{i}=\left[\begin{array}[]{ccc}0&\frac{1}{C_{t,i}}&0\\ \frac{k_{1,i}-1}{L_{t,i}}&\frac{k_{2,i}-R_{t,i}}{L_{t,i}}&\frac{k_{3,i}}{L_{t,i}}\\ -1&0&0\end{array}\right],~B_{i}=\left[\begin{array}[]{c}0\\ 0\\ 1\end{array}\right],~D_{i}=\left[\begin{array}[]{c}-\frac{1}{C_{t,i}}\\ 0\\ 0\end{array}\right],~E_{i}=\begin{bmatrix}0\\ \frac{1}{L_{t,i}}\\ 0\end{bmatrix},~C_{i}={\bf I}_{3}.\end{split}$
Since the voltage $V_{i}$ and the filter current $I_{t,i}$ are measurable [26], and $v_{i}$ is the controller variable, we have full state measurement here, i.e., $C_{i}={\bf I}_{3}$. Let us further introduce the following assumption on noise.
@@
The covariance matrices $\Sigma_{\delta_{i}}$, $\Sigma_{\epsilon_{k}}$ and $\Sigma_{\zeta_{i}}$ can be computed from available information about the local models, sensors, and possibly on historical data through Monte Carlo-based approaches [27], and a discussion of their computational approach can be found in [28]. In addition, we account for unmodeled system uncertainties within the stochastic process noise, as commonly adopted in literature [29, Section 3.4].
@@
This work aims to develop a distributed method to detect and estimate both the actuator fault $f_{a,i}$ and the aggregated faulty line current $f_{I,i}$ in DG unit $i$, while accounting for the unknown power load $P_{i}$ and the noise terms $\delta_{i}$$\epsilon_{k}$, and $\zeta_{i}$. Estimating $f_{a,i}$ is relatively straightforward, as it is the only unknown signal in the dynamics of the measurable current $I_{t,i}$ in (1). However, it is challenging to even detect $f_{I,i}$ because the line current $I_{k}$ is unavailable and $f_{I,i}$ is coupled with $P_{i}$ in $d_{i}$, as described in (4). It is worth noting that $f_{I,i}$ and $P_{i}$ are indistinguishable in the following two cases:
@@
    Arbitrarily varying $P_{i}$ and $f_{I,i}$. If $P_{i}(t)$ and $f_{I,i}(t)$ can take any values, different combinations of $P_{i}(t)$ and $f_{I,i}(t)$ can yield identical total values and result in an indistinguishable scenario. For instance, consider two different pairs, $(P_{i}(t),f_{I,i}(t))$ and $(P^{\prime}_{i}(t),f^{\prime}_{I,i}(t))$. It holds that
@@
    Constant $V_{i}$. If $V_{i}(t)$ is constant, separation of $f_{I,i}$ and $P_{i}$ is theoretically impossible because the signal $V_{i}$ does not provide sufficient variation to distinguish their contributions in the aggregated signal $P_{i}/V_{i}+f_{I,i}$. This is close to the well-known persistence of excitation phenomena for LTI systems, as discussed in [30, Theorem 3.5]. It is also reflected in the rank condition following (30) in Subsection 4.4.
Fortunately, in practical DC microgrids, power loads do not vary arbitrarily but typically follow specific patterns. Various types of loads, such as data centers, electric vehicle charging stations, motor drives, and inverter-based AC loads, can be approximated as constant power loads [31, 1]. Based on this observation, the following assumption about $P_{i}$ is made throughout the paper to exclude the first scenario where $P_{i}$ and $f_{I,i}$ become indistinguishable. We emphasize that no assumption is imposed on the fault signals.
@@
The power load $P_{i}(t)$ is a piece-wise constant signal, with each value persisting for a sufficiently large duration.
In the second indistinguishable case, although both faults and load changes can cause variations in $V_{i}$, these effects are rapidly attenuated by feedback control, particularly in scenarios involving incipient faults or step load changes. As a result, $V_{i}$ typically exhibits slow variations or converges to a constant value, rendering the estimation problem of $f_{I,i}$ from $P_{i}/V_{i}+f_{I,i}$ ill-posed.
@@
    where $\bar{f}_{I,i}(t)$ represents the average value of $f_{I,i}(t)$ over a period, $\mathcal{C}\left(\cdot\right)$ is a time-varying bound dependent on the system and power line dynamics (2)-(4), input signals to $\mathcal{D}_{l,i}$, and properties of unknown signals $P_{i}$ and $f_{I,i}$.

### v2

@@ Abstract.
This paper proposes a distributed diagnosis scheme to detect and estimate actuator and power line faults in DC microgrids (e.g., electric-vehicle charging microgrids) subject to unknown power loads and stochastic noise. To address actuator faults, we develop an optimization-based filter design approach within the differential-algebraic equation (DAE) framework, which achieves fault estimation, decoupling from power line faults, and robustness against noise. In contrast, the estimation of power line faults poses greater challenges due to the inherent coupling between fault currents and unknown power loads, especially under insufficient system excitation, where their effects become difficult to distinguish from measurements. To the best of our knowledge, this is the first study to address this critical yet underexplored issue. Our solution introduces a novel differentiate-before-estimate strategy. A set of diagnosis rules based on the temporal characteristics (i.e., duration of threshold violation) of a constructed residual is developed to distinguish step load changes from line faults. Once a power line fault is detected, a regularized least-squares (LS) method is activated to estimate the fault currents, for which we further derive an upper bound on the estimation error. Finally, comprehensive simulations validate the effectiveness of the proposed scheme in terms of estimation accuracy and robustness against disturbances and noise under different fault scenarios.
@@
Jingwei Dong (jingwei.dong@it.uu.se), Per Mattsson, and André Teixeira are with the Division of Systems and Control, Uppsala University, Sweden. Mahdieh S. Sadabadi is with the Department of Electrical and Electronic Engineering, The University of Manchester, Manchester, U.K.
@@ 1.1. Literature Review
The existing results on fault diagnosis of DC microgrids typically focus on fault detection and isolation. The most straightforward way is to set detection thresholds based on characteristics of measured currents and voltages, as presented in [9]. While thresholding-based methods are easy to implement, their diagnosis performance is highly sensitive to the selection of thresholds. By exploring additional time-domain information from measured signals, methods based on differential currents [4, 15] and traveling wave analysis [25] can achieve more reliable diagnosis results compared to simple thresholding-based approaches. For instance, the authors in [15] analyzed the first and second derivatives of fault currents to detect DC cable ground faults in the presence of load and operating mode changes. However, such derivative-based approaches are sensitive to noise. In addition to time-domain information, frequency-domain methods, such as fast Fourier transform [28] and wavelet transform [38], are also effective for fault analysis in microgrid systems. Yet, their reliance on case-specific fault frequency characteristics limits their generalization capability. In recent years, the availability of large amounts of historical data has also facilitated the adoption of learning-based approaches in DC microgrid fault diagnosis. A key challenge, however, is the limited amount of data in fault scenarios. We refer interested readers to [21] and [37] for further insights.
@@
Let us first briefly review some observer-based fault diagnosis methods developed for DC microgrids. In [39], an unknown input observer was employed to detect arc faults on power lines. To enhance robustness against disturbances while maintaining fault sensitivity, the authors in [35] utilized the mixed $\mathcal{H}_{\_}/\mathcal{H}_{\infty}$ index in the design of an optimal fault detection observer for DC microgrids. Fault isolation can be achieved by building a bank of observers, each of which is designed to be sensitive to specific faults. For instance, the authors in [36] constructed fault matrices (or signatures) for different types of faults in DC microgrids, and then designed a bank of unknown input observers based on corresponding sensitivity matrices to distinguish fault types. In [31], a bank of unscented Kalman filters was developed to estimate system states under faults, whose outputs were then integrated to locate faulty sensors in DC microgrids. Regarding DAE-based approaches, they offer several advantages over conventional observer-based methods (e.g., Kalman filters, sliding mode observers, and adaptive observers), including: (i) characterizing all possible residual generators for systems described by DAEs, (ii) deriving residual generators of the lowest possible order [19], (iii) facilitating the incorporation of multiple design objectives and constraints, with the corresponding parameters determined within a well-defined optimization framework [16]. These benefits have also led to the development of anomaly detection methods for microgrid systems within the DAE framework, such as attack detection in [20] and ground fault detection in [7].
@@
Note that the aforementioned methods primarily focus on fault detection and isolation, while research on fault estimation for DC microgrids remains limited. Fault estimation involves determining the shape and size of faults, which has more stringent requirements than fault detection and isolation. The complexity is further exacerbated in DC microgrids with unknown power loads due to the introduction of nonlinearity [1]. To address this issue, the authors in [1] employed a sliding mode observer to estimate actuator and sensor faults in DC microgrids. More recently, an adaptive fault estimation method was developed in [33] to estimate the system states and faults simultaneously. However, both [1] and [33] rely on power line or load current measurements, which require additional sensors that increase system complexity and maintenance costs. The authors in [3] addressed this limitation by proposing a method based on an augmented state observer and an open-loop line current estimator, but the method is confined to sensor attacks.
@@
Based on the above analysis, the problem of multiple fault detection and estimation that simultaneously considers unmeasurable line currents and unknown power loads in DC microgrids remains insufficiently explored. The gap is reflected in two main aspects. First, most existing signal analysis-based [9, 4, 15, 25] and model-based methods [1, 33] are developed under the assumption that power line currents at one or multiple ends are measurable, which requires additional sensing and communication infrastructure and thus increases deployment cost. When considering unmeasurable line currents, power line faults are no longer directly observable but must instead be inferred indirectly from the system dynamics. Second, research on unknown power loads in DC microgrids has mainly focused on system stability analysis and control design, such as those in [12, 23]. In the context of fault detection and estimation, most existing studies simplify the problem by considering only impedance-type loads or assuming measurable load currents, which substantially reduces the difficulty of analysis. We further observe that the coupled effects of fault-induced line currents and unknown power loads pose great challenges in distinguishing between them. This difficulty is exacerbated under insufficient system excitation (e.g., in the case of incipient faults), where the slow variation of signals leads to an ill-posed estimation problem. To the best of our knowledge, this is the first study to address this challenge in the context of DC microgrids.
@@ 1.2. Main Contributions
  • Multiple fault estimation framework. We develop a multiple fault estimation framework for DC microgrids, which accounts for both actuator and power line faults in the presence of unknown power loads and stochastic noise. This framework is built on fault estimation filters developed within the DAE framework, where the filter design is formulated as a tractable optimization problem (Proposition 3.1).
  • Differentiation strategy for power line faults and step power load changes. Building on the practical assumption of piecewise-constant power loads, we propose a novel strategy to distinguish between power line faults and step load changes by exploiting the temporal characteristic differences of their induced residuals, specifically, the duration of threshold violation (Proposition 4.1). This differentiation process is further formalized into a set of diagnosis rules (Equation (27)), facilitating the subsequent faulty line current estimation.
  • Faulty line current estimation with an explicit error bound. To address the ill-posed issue in faulty line current estimation, we formulate a regularized least-squares (LS) optimization method. An analytical solution is further derived to enable real-time estimation (Proposition 4.3). Furthermore, we establish a computable bound on the expectation of the estimation error, explicitly characterizing its dependence on load and fault variations (Theorem 4.4).
@@ 2.1. DG Dynamics
where $V_{i}(t)\in{\mathbb{R}}$ and $I_{t,i}(t)\in{\mathbb{R}}$ are the voltage at the Point of Common Coupling (PCC) $i$ and the filter current, respectively. The unknown power demand is denoted by $P_{i}(t)\in{\mathbb{R}}_{+}$. The control signal $u_{i}(t)=\varphi_{i}(t)V_{dc,i}$, where $\varphi_{i}(t)$ is the duty cycle of the buck converter. The current of power line $k$ is represented by $I_{k}(t)\in{\mathbb{R}}$ where $k\in\mathbb{N}_{L}$. The microgrid topology is encoded in the incidence matrix element $\mathbb{B}_{ik}$, which indicates the connection between power line $k$ and DG unit $i$. In particular, $\mathbb{B}_{ik}=1$ if DG unit $i$ is the positive end of line $k$, $\mathbb{B}_{ik}=-1$ if DG unit $i$ is the negative end of line $k$, and $\mathbb{B}_{ik}=0$ otherwise. Therefore, the term $\sum_{k=1}^{m}\mathbb{B}_{ik}I_{k}(t)$ is the total current injected into DG unit $i$ from its neighbors and represents the physical couplings of DG unit $i$ with neighboring DG units. The additive actuator fault in DG unit $i$ is denoted by $f_{a,i}(t)\in{\mathbb{R}}$.
@@ 2.2. Power Line Dynamics
where $R_{k}$ and $L_{k}$ denote the line resistance and inductance, respectively. The fault on line $k$ is denoted by $f_{L,k}(t)\in{\mathbb{R}}$.
@@ 2.3. Stabilizing Voltage Controller
where $V_{i}^{*}$ is a reference voltage provided by a higher-level controller, $K_{i}=[k_{1,i}\ k_{2,i}\ k_{3,i}]$ is the voltage control gain vector. Design methods for $K_{i}$ can be found in [23, 29].
@@ 2.4. Closed-loop Dynamics
where $f_{I,{i}}(t)=\sum_{k=1}^{m}\mathbb{B}_{ik}I_{k,f}(t)$ denotes the aggregated faulty line current affecting DG unit $i$. The system matrices in (4) are as follows:
$\begin{split}&A_{i}=\left[\begin{array}[]{ccc}0&\frac{1}{C_{t,i}}&0\\ \frac{k_{1,i}-1}{L_{t,i}}&\frac{k_{2,i}-R_{t,i}}{L_{t,i}}&\frac{k_{3,i}}{L_{t,i}}\\ -1&0&0\end{array}\right],~B_{i}=\left[\begin{array}[]{c}0\\ 0\\ 1\end{array}\right],~D_{i}=\left[\begin{array}[]{c}-\frac{1}{C_{t,i}}\\ 0\\ 0\end{array}\right],~E_{i}=\begin{bmatrix}0\\ \frac{1}{L_{t,i}}\\ 0\end{bmatrix},~C_{i}={\bf I}_{3}.\end{split}$
Since the voltage $V_{i}$ and the filter current $I_{t,i}$ are measurable [30], and $v_{i}$ is the controller variable, we have full state measurement. Let us further introduce the following assumption on noise.
@@
The covariance matrices $\Sigma_{\delta_{i}}$, $\Sigma_{\epsilon_{k}}$ and $\Sigma_{\zeta_{i}}$ can be computed from available information about the local models, sensors, and possibly on historical data through Monte Carlo-based approaches [2], and a discussion of their computational approach can be found in [22]. In addition, we account for unmodeled system uncertainties within the stochastic process noise, as commonly adopted in literature [5, Section 3.4].
@@
This work aims to develop a distributed method to detect and estimate both the actuator fault $f_{a,i}$ and the aggregated faulty line current $f_{I,i}$ in DG unit $i$, while accounting for the unknown power load $P_{i}$ and the noise terms $\delta_{i}$$\epsilon_{k}$, and $\zeta_{i}$. Estimating $f_{a,i}$ is relatively straightforward, as it is the only unknown signal in the dynamics of the measurable current $I_{t,i}$ in (1) except for noise. However, it is challenging to even detect $f_{I,i}$ because the line current $I_{k}$ is unavailable and $f_{I,i}$ is coupled with $P_{i}$ in $d_{i}$, as described in (4). From an algebraic perspective, $f_{I,i}$ and $P_{i}$ are fundamentally indistinguishable under the following two scenarios:
@@
    Arbitrarily varying $P_{i}$ and $f_{I,i}$. If $P_{i}(t)$ and $f_{I,i}(t)$ are unconstrained, different combinations of $P_{i}(t)$ and $f_{I,i}(t)$ can yield identical effects. For instance, for two different pairs $(P_{i}(t),f_{I,i}(t))$ and $(P^{\prime}_{i}(t),f^{\prime}_{I,i}(t))$, it holds that
@@
    Constant $V_{i}$. If $V_{i}(t)$ is constant, separation of $f_{I,i}$ and $P_{i}$ is theoretically impossible because $V_{i}$ provides no variation that can be exploited to distinguish them from the aggregated signal $P_{i}/V_{i}+f_{I,i}$. This is close to the well-known persistence of excitation phenomena for LTI systems, as discussed in [32, Theorem 3.5]. It is also reflected in the rank condition following (28) in Subsection 4.4.
Beyond the above algebraic conditions, the indistinguishability between $P_{i}$ and $f_{I,i}$ also has a physical origin in DC microgrid dynamics. As shown in Fig. 1, both load variations and line faults act as current-type disturbances injected at the same point, and therefore share identical spatial signatures. Moreover, the voltage regulation loop suppresses voltage variations, resulting in insufficient excitation in the measured signals and further weakening the separability of $P_{i}$ and $f_{I,i}$. Fortunately, in practical DC microgrids, $V_{i}$ does vary due to faults and load changes, and power loads do not evolve arbitrarily but typically follow specific patterns. Motivated by this observation, we impose the following assumption on $P_{i}$ throughout the paper.
@@
The power load $P_{i}$ is a piece-wise constant signal that allows multiple step changes over time, with each value persisting for a sufficiently large duration.
Note that assuming constant or piecewise-constant power loads is a standard modeling practice in both control design and fault diagnosis for DC microgrids (e.g., [11, 23]). This assumption is also practically motivated, as many real-world loads, such as data centers, electric-vehicle charging stations, motor drives, and inverter-based AC loads [12, 17], can be reasonably characterized as instantaneous constant power loads. We emphasize that no assumption is imposed on the fault signals.
@@
    where $\bar{f}_{I,i}(t)$ represents the average value of $f_{I,i}$ over a period, $\mathcal{C}\left(\cdot\right)$ is a time-varying bound dependent on the system and power line dynamics (2)-(4), input signals to $\mathcal{D}_{l,i}$, and properties of unknown signals $P_{i}$ and $f_{I,i}$.